OpenStudy (anonymous):

NO ONE IS HELPING TODAY LIKE I HAVE ASKED 50 TIMES AND GOT NOTHIN I NEED HELP

OpenStudy (anonymous):

y = 7x − 8 y = 5x − 2 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. (5 points) Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer. (5 points)

OpenStudy (anonymous):

IT HAS TO B IN ELIMINATION OR SUBSTITION

OpenStudy (anonymous):

srry caps was on

OpenStudy (anonymous):

@Wolf_Girl2002

OpenStudy (anonymous):

@MichelleiSMine

OpenStudy (anonymous):

@curiositygirl

OpenStudy (anonymous):

btw i provide medal

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

i did this last year so hold on while I try to refresh my memory on it

jabez177 (jabez177):

WOW I CAN TRY AND DO MY BEST BUT IT IS A BIT 2 HARD

OpenStudy (anonymous):

ok thx guyz u da best

OpenStudy (anonymous):

and sadly this type of equation was my fav and i just forgot how to do it XD

jabez177 (jabez177):

YEAH WHAT SHE SAID SORRY

OpenStudy (anonymous):

awe man...XD well i need it really bad i horible at math like really bad

jabez177 (jabez177):

@austin11111 CAN U HELP

jabez177 (jabez177):

@ashhhrodriguez CAN U HELP PLZ

OpenStudy (anonymous):

thx...@austin11111

OpenStudy (anonymous):

@austin11111

jabez177 (jabez177):

@iambatman @mathmath333 CAN U HELP

OpenStudy (anonymous):

i may be able to help

OpenStudy (anonymous):

thx u so much

OpenStudy (anonymous):

but it has to b in elimination or substitution

OpenStudy (anonymous):

Sure, I can help you

OpenStudy (anonymous):

yay thx u guyz

OpenStudy (anonymous):

um sorry guys im not the best in mathematics

OpenStudy (anonymous):

o ok np

OpenStudy (anonymous):

Alright what method do you want to use?

OpenStudy (anonymous):

Substitution or elimination?

OpenStudy (anonymous):

substitution

jabez177 (jabez177):

THNX BAT 4 HELPING HIM

OpenStudy (anonymous):

The idea for substitution is, you take one equation, isolate one of the variables, then plug it into the other equation, so for example: You have, y = 7x − 8 y = 5x − 2 Well, they already made it simple for you, as they gave you solved variables (for y). Now let y = y \[5x-2=7x-8\] solve for x here. Once you get x, that will be one point to a solution, the next step is to solve for y, this is even easier because you take your x value and plug it in for either of the equations, take the first one or the second one.

OpenStudy (anonymous):

are u asking me to take the first one or the second one i would pic the first

OpenStudy (anonymous):

Well first solve for x, then plug it in either of the equations to get a corresponding y, it doesn't matter which one you pick.

OpenStudy (anonymous):

When you get the (x,y) that will be the point where the lines intersect.

OpenStudy (anonymous):

i usually went for the one I thought was easier

OpenStudy (anonymous):

so what is the value of x im sorry im really not good with math

OpenStudy (anonymous):

Ok so you have two equations right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

set y = y 5x-2=7x-8 and solve for x.

OpenStudy (anonymous):

so it would be 3x and 1x or idk

OpenStudy (anonymous):

How did you get that? I see there is a problem you're having with algebra.

OpenStudy (anonymous):

i subtracted each side and yea i am havin a problem with it it was never my strong subject

OpenStudy (anonymous):

did you subtract the -2 from the -8?

OpenStudy (anonymous):

no 2 from the 5 and 7 from the 8 they not negative the subtracting

OpenStudy (anonymous):

Generally when you have an equation, what you want to do is, isolate a variable meaning you want that variable on its own, and what you do to the left side you must also do to the right, so for example lets take: x+3=3x-3 and I ask you to solve for x. So lets move all the constants on one side and the variables on the other, so I'm going to move the x's to the left and the 3's to the right in doing so I use simple operations that we all know, add, subtract, multiply, divide. Lets first move the x's: x-3x+3=3x-3-3x you see the 3x on the right side gets cancelled out so it balances the left. So now we have x-3x+3 = -3 Let's now move the 3 to right side. What's the opposite operation of adding? Subtracting right. x-3x+3-3=-3-3 x-3x=-6 We're almost done, so now we have: -4x=-6, ok so the -4 is "stuck" onto the x, meaning it's multiplied, what's the opposite of multiplying? Dividing. Divide -4 on both sides, x=-6/-4 = 6/4 = 3/2

OpenStudy (anonymous):

Now use this and try to solve for x :)

OpenStudy (anonymous):

ohhh thats what you did. your supposed to try to get the variable on one side. try adding the -2 and the -8 to get the 5x by itself thats what i was taught last year

OpenStudy (anonymous):

Sorry for using so many 3's in my example haha...it's early >.<

OpenStudy (anonymous):

ikr!

OpenStudy (anonymous):

np...thx i think i get it

OpenStudy (anonymous):

I will be back much later, I have to go to class now, take care and good luck and I see you're good hands with curiositygirl. That's great @srgntb0$$ :)

OpenStudy (anonymous):

can u stay here so i can do the problem and c if i get it right

OpenStudy (anonymous):

I would love to srgnt, but work on it, and I will check in with you later, I'm getting late, so I'll catch you later!

OpenStudy (anonymous):

o ok cya

OpenStudy (anonymous):

have fun in class iambatman

OpenStudy (anonymous):

r u stayin

jabez177 (jabez177):

I GUES BATMAN IS FRIENDLY XD

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

OMG I CANT DO IT I DONT UNDERSTAND IT ON MY PROBLEM

OpenStudy (anonymous):

okay hold on. there is like nothing to freak out about trust me. when i first learned how to do it i was in the same boat as you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it fine i just gona try to get everything else right on it and if i get this wrong ill at least get an 80%...what school u in i in flvs

OpenStudy (anonymous):

im in flvs also. and what I did so far is set it up like iambatman said to do 5x-2=7x-8 I then added the -2 to the -8 and got -6

OpenStudy (anonymous):

where did u get the negative 8 and 2 they were subtracting from the 5x and 7x...srry it takin so long i suck at math

OpenStudy (anonymous):

well since the equation is y=7x-8 y=5x-2 for substitution you set it up like this> 5x-2=7x-8

OpenStudy (anonymous):

and since you need to get the x by itself you want to make either 5x be by itself or 7x

OpenStudy (anonymous):

ok but idk how to find the answer and if u explain it i wont understand i didnt understand imbatman's

OpenStudy (anonymous):

u in 8th arent u i in 7th

OpenStudy (anonymous):

lol its okay i am actually trying to make it as simple as i had actually im in 9th

OpenStudy (anonymous):

o cool...if i get the 80% at least i have a b that all i need to pass but i still need something resonable to put on the written part

OpenStudy (anonymous):

yeah thatll be good do you want me to give you part of the written part?

OpenStudy (anonymous):

sure...also i need the answer to another question

OpenStudy (anonymous):

Two lines, C and D, are represented by the equations given below: Line C: y = x + 14 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? (6, 20), because both lines pass through this point (6, 20), because the point does not lie on any axis (3, 11), because one of the lines passes through this point (3, 11), because the point lies between the two axes

OpenStudy (anonymous):

Wilson is 6 years older than Mary. Wilson's age is 10 years less than five times Mary's age. The equations below model the relationship between Wilson's age (w) and Mary's age (m): w = m + 6 w = 5m − 10 Which is a possible correct method to find Wilson's and Mary's ages? Solve w + 6 = 5w − 10 to find the value of w. Write the points where the graphs of the equations intersect the x axis. Solve m + 6 = 5m − 10 to find the value of m. Write the points where the graphs of the equations intersect the y axis.

OpenStudy (anonymous):

The work of a student to solve the equation 3(2x − 4) = 8 + 2x + 6 is shown below: Step 1: 3(2x − 4) = 8 + 2x + 6 Step 2: 5x − 7 = 14 + 2x Step 3: 5x − 2x = 14 + 7 Step 4: 3x = 21 Step 5: x = 7 In which step did the student first make an error and what is the correct step? Step 2; 6x − 12 = 14 + 2x Step 2; 6x − 7 = 2(6 + x + 4) Step 3; 5x − 2x = 14 − 7 Step 3; 5x + 2x = 14 + 7

OpenStudy (anonymous):

@curiositygirl i thought u were gona help me

OpenStudy (anonymous):

okay For substitution you want to get the x by itself. so you set up the equation by eliminating y. you are then left with 5x-2=7x-8. to get the x by itself you want to move the -2 over and combine it with the -8. this becomes 5x=7x-6 after adding the -2 with the -8 sorry if its confusing

OpenStudy (anonymous):

so is that the answer srry kinda confusing...and what about the others and y dont u have a profile pic

OpenStudy (anonymous):

thats part of Part A because idk how to explain the rest i never thought of having one yet

OpenStudy (anonymous):

hold on about the others i am trying to figure them out

OpenStudy (anonymous):

ok thx u y dont u got a profile pic

OpenStudy (anonymous):

welcome idk

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

hay i think y=13 and x=3 is that right

OpenStudy (anonymous):

idk what

OpenStudy (anonymous):

about the other questions what are they

OpenStudy (anonymous):

@sashadqaf can u plz help me

OpenStudy (anonymous):

id help but im really really really bad at math :c sowy

OpenStudy (anonymous):

@KamiBug plz help

OpenStudy (anonymous):

can u help

OpenStudy (kamibug):

Which question is it?

OpenStudy (anonymous):

the last 4 i asked

OpenStudy (kamibug):

Okay, give me sec. :)

OpenStudy (anonymous):

ok i go to thesame school as u flvs

OpenStudy (anonymous):

@philcua can u help

OpenStudy (anonymous):

i need a lot